OpenStudy (anonymous):

Use Cramer's Rule to find solution set: 4x-y=2 3x+2y=7

OpenStudy (turingtest):

do you know how to find the determinant of a matrix?

OpenStudy (turingtest):

@mathsux4real please engage those who are trying to help you

OpenStudy (anonymous):

Left the computer. My fault. And no I do not.

OpenStudy (turingtest):

a determinant of a 2x2 matrix is given by\[\left|\begin{matrix}a&b\\c&d\end{matrix}\right|=ad-cb\]

OpenStudy (anonymous):

let \[b= \left(\begin{matrix}2 \\ 7\end{matrix}\right)\] \[A=\left[\begin{matrix}4 & -1 \\ 3 & 2\end{matrix}\right]\] then by cramer's rule: \[x=\det A_1/detA \] \[y=detA_2/detA\] where A_i is the matrix obtained by replacing the ith column of A with b. first let's find detA_1,detA_2, and det A \[detA=\det \left[\begin{matrix}4 & -1 \\ 3 & 2\end{matrix}\right]=(4)(2)-(-1)(3)=8+3=11\] \[detA_1=\det \left[\begin{matrix}2 & -1 \\ 7 & 2\end{matrix}\right]=(2)(2)-(-1)(7)=11\] \[detA_2=\det \left[\begin{matrix}4 & 2 \\ 3 & 7\end{matrix}\right]=(4)(7)-(3)(2)=22\] x=detA_1/detA=11/11=1 y=detA_2/detA=22/11=2 x=1 y=2

OpenStudy (anonymous):

Wow anonymoustwo44, thank you so much! I really appreciate the thorough explanation. Really helps

OpenStudy (anonymous):

yw!

OpenStudy (anonymous):

It sucks that you can't give a medal to more than one person :/